can some one help me find the diameter.
Diameter of a circle
If the diameter of a circle is 1.3 * 10^-12 meters, then what is the radius.

Answer by stanbon(75887)

Diameter of a circle
If the diameter of a circle is 1.3 * 10^-12 meters, then what is the radius.
-------------
radius = (1/2)diameter
radius = (1/2)*(1.3)*10^-12
=0.65*1*10^-12
=6.5*10^-13 meters
Cheers,
Stan H.